# CP Violation of Baryon Decays from $N\pi$ Scatterings

### v1

@@ Abstract
It can be seen from the above formula that the interference effects in the tree and penguin contributions are different. The above difference comes from the signs of the $a_{6}$ terms, compared to those of the $a_{1}$ and $a_{4}$ terms. In the four-fermion operators of $O_{1-4}$, quarks are all left-handed due to the weak interaction $(\bar{u}b)_{V-A}=\bar{u}_{L}\gamma^{\mu}b_{L}$. On the contrary, the $u$ quark from the $b$ transitions is right-handed with the operators of $O_{5,6}$, $(\bar{u}b)_{S-P}=\bar{u}_{R}b_{L}$. The different chiralities of $u$ quark result in different helicities of the $N\pi$ system, and hence different interference effects of partial-wave amplitudes. In the end, they contribute to the difference of strong phases for CPV.
@@
Benefited by the partial-wave amplitudes, it can be easily studied for the angular analysis on $\Lambda_{b}$ decays. The angular distributions are illustrated in Fig. 2 for $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The details of angular distributions can be easily obtained by the helicity amplitudes. The angle $\theta$ is defined between the directions of momentum of the final baryon $p(\Delta^{++})$ in the rest frame of $N\pi$ system and the momentum of $N\pi$ in the rest frame of $\Lambda_{b}^{0}$.
Numerical Discussions.
Table 1: Global CPV of $\Lambda_{b}$ decays stemming from $N\pi$ scatterings.
[…]
$N\pi\to\Delta^{++}\pi^{-}$ $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$ 6.0
$m_{N\pi}\in[1.2,1.9]$GeV $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})\pi^{-}$-4.4
$N\pi\to p\pi^{0}$ $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$ 5.8
$m_{N\pi}\in[1.1,2.5]$GeV $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$-3.9
@@
Based on the above CPV mechanism and the data of $N\pi$ scatterings, we can predict the global and regional CP asymmetries of $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The inputs of effective Wilson coefficients are given in Ali:1998eb . The inputs of $N\pi$ scattering amplitudes are model-independent and given in SAID . Since the data of scatterings of $N\pi\to N\pi$ and $\Delta\pi$ are very precise, we neglect the uncertainties of scattering amplitudes. Little is known about the form factors of $\Lambda_{b}\to N\pi$ at the current stage. But it is similar to the transition form factors of $B\to\pi\pi$ which have been studied by the heavy meson chiral perturbation theory Lee:1992ih ; Fajfer:1998yc ; Cheng:2007si ; Cheng:2020ipp and the light-cone QCD sum rules Cheng:2017smj ; Cheng:2017sfk . Therefore the form factors of $\Lambda_{b}\to N\pi$ can be calculated using the proven methods in principle. At the current stage, we will take some simple and typical values of these form factors, in order to reveal the new CPV mechanism. The point here is to show that large regional CPV can be generated in this mechanism, but not do a very precise prediction. Besides, the absolute values of form factors are cancelled in the ratio of CPV definition. So CPV relies only on the relative values of form factors. In the heavy quark limit, $f_{1}^{i}(0)=g_{1}^{i}(0)$. It is known that the heavy quark symmetry violating effect in the heavy-light baryonic form factors could be about $20\%$ Han:2022srw ; Huang:2022lfr . Therefore, we will consider a heavy-quark symmetry violating effect as $20\%$ without loss of generality. Besides, the mass regions of $p\pi^{0}$ and $\Delta^{++}\pi^{-}$ are taken as what SAID program provides, i.e. $m_{p\pi^{0}}\in[1.1,2.5]$GeV and $m_{\Delta^{++}\pi^{-}}\in[1.2,1.9]$GeV.
We can get the numerical results using the above inputs. Firstly, the global CPV in the mass regions of the $N\pi$ scattering system are given in Table. 1. It can be found that the magnitudes of global CPV of these processes are around $(4-6)\%$. Such values of CPV could be measured by LHCb for $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ and $(p\pi^{+}\pi^{-})\pi^{-}$ whose data samples are large currently with all the charged final-state particles.
Figure 3: CPV distributions of the Dalitz plots. The four plots are $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$ and $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, respectively.
@@
The regional CPV in the Dalitz plots are even more interesting. Every partial-wave amplitude corresponds to different angular distribution. Then the interference effects in the different regions in the Dalitz plots would induce different CP asymmetries. The numerical results of regional CPV are given in Fig. 3. It could found that the CPV could be as large as $20\%$ or even larger in some regions. This is large enough for the observation in experiments.
@@
One more interesting property in the Dalitz plot is that there exists large regions with CPV of positive or negative values. This would be helpful for the experimental measurements. For example, it can be divided into two bins with $\cos\theta>0$ and $\cos\theta<0$, or say roughly $m_{hh}>1.5$GeV and $m_{hh}<1.5$GeV with $h=\pi$ or $K$, in all the four processes. The CPV in each bin could be large enough for experimental measurements, with large enough data samples.
@@
From the Dalitz plots, it could be clearly seen that the regional CPV depend on $\cos\theta$ significantly. This is reasonable that the interference betwee different partial-wave amplitudes induce the dependence of CPV on $\cos\theta$. In order to see the dependence, it is very convenient to express the differential decay width on $\cos\theta$. The Fourier distributions after integrating out scattering energy are
$\frac{d\Gamma}{d\cos\theta}\propto\sum^{n=5}_{n=0}\mathcal{C}_{n}\cos n\theta$ (3)
with $\mathcal{C}_{n}=(1,-0.12,0.17,-0.02,0.003,0.02)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, $(1,-0.29,0.14,-0.04,-0.05,0.02)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$, (1,1.3,0.7,0.8,0.03,0.13,(1,-1.3,0.7,-0.8,-0.03,-0.13,( 1 , - 1.3 , 0.7 , - 0.8 , - 0.03 , - 0.13 , 0.004)-0.004)- 0.004 ) for $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$, and (1,4.3,2.6,1.3,(1,-4.3,2.6,-1.3,( 1 , - 4.3 , 2.6 , - 1.3 , 0.4,0.2,0.02)0.4,-0.2,0.02)0.4 , - 0.2 , 0.02 ) for $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, respectively. It can be found that the first two moments of $\cos\theta$ are large, so that the CPV in the Dalitz plots are basically caused by $\cos\theta$.
Figure 4: CPV of Fourier moments of $\cos(n\theta)$. The four plots are $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$ and $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, respectively.
@@
So then the large regional CPV could also be measured by the moments of $\cos(n\theta)$. It can be easily realized by the experiments using the differential width convoluted with the moments of $\cos(n\theta)$. The results of CPV of Fourier moments are given in Fig. 4. The interesting result is that the CPV of the first moment $\cos\theta$ could be as large as $40\%$ in $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$. Considering the large component of the $\cos\theta$ term in the Eq. (3) for $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$, it has a large possibility to observe the CPV of the $\cos\theta$ moment.
@@
In the end, we strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ in the mass region of $m_{p\pi\pi}\in[1.2,1,9]$GeV, considering that $\Lambda_{b}^{0}\to p\pi^{+}\pi^{-}K^{-}$ has the largest data sample of charmless $b$-baryon decays. It could be in three ways. Firstly, the global CPV with the value of $6\%$ could be reached with all the data sample in the considered mass region. Secondly, the CPV in the Dalitz plot can be measured by two bins of $\cos\theta>0$ and $\cos\theta<0$, with the regional CPV of $(10-20)\%$. Thirdly, it is also realizable to measure the CPV of moment of $\cos\theta$ with a value as large as $40\%$.

### v3

@@ II CPV mechanism with $N\pi$ scatterings
It can be seen from the above formula that the interference effects in the tree and penguin contributions are different. The above difference comes from the signs of the $a_{6}$ terms, compared to those of the $a_{1}$ and $a_{4}$ terms. In the four-fermion operators of $O_{1-4}$, quarks are all left-handed due to the weak interaction $(\bar{u}b)_{V-A}=\bar{u}_{L}\gamma^{\mu}b_{L}$. On the contrary, the $u$ quark from the $b$ transitions is right-handed with the operators of $O_{5,6}$ after the Feirz transformation, $(\bar{u}b)_{S-P}=\bar{u}_{R}b_{L}$. The different chirality of the $u$ quark results in different helicity of the $N\pi$ system in the relativistic case, and hence different interference effects of partial-wave amplitudes.
There are two sources of difference of strong phases. One is from the effective Wilson coefficients of $a_{4}$ and $a_{6}$ where the quark loops, the chromo-meganetic operator, and the vertex corrections could contribute the strong phases in the generalized factorization hypothesis Ali:1998eb . It basically contributes to the global CPV. The second source of difference of strong phases is from the different interfering effects between the partial-wave amplitudes of $N\pi$ scatterings. This mainly leads to the regional CPV whose values might be significantly large in some regions of the Dalitz plot where the difference of strong phases could be very large.
The two sources of CP violation can be clearly illustrated through an analytical example involving two partial-wave amplitudes $S_{11}$ and $P_{11}$, whereas CPV with higher partial waves are very similar. The partial width with only these terms is given by
$d\Gamma$ $\propto|P_{11}|^{2}(|A|^{2}+\kappa^{2}|B|^{2})+|S_{11}|^{2}(|\tilde{A}|^{2}+\kappa^{2}|\tilde{B}|^{2})$ (5)
$+2\mathcal{R}e\left[(A\tilde{A}^{*}+\kappa^{2}B\tilde{B}^{*})P_{11}S^{*}_{11}\right]$
with
$A$ $=(\lambda_{u}a_{1}-\lambda_{t}a_{46+})f^{\frac{1}{2}+}_{1}m_{-}$ (6)
$B$ $=(\lambda_{u}a_{1}-\lambda_{t}a_{46-})g^{\frac{1}{2}+}_{1}m_{+}$
$\tilde{A}$ $=(-\lambda_{u}a_{1}+\lambda_{t}a_{46-})f^{\frac{1}{2}-}_{1}m_{-}$
$\tilde{B}$ $=(-\lambda_{u}a_{1}+\lambda_{t}a_{46+})g^{\frac{1}{2}-}_{1}m_{+}$
which can be derived directly from Eq.(4). The parameter $\kappa=p_{N\pi}/(E_{N\pi}+m_{N\pi})$, with $p_{N\pi}$, $E_{N\pi}$ and $m_{N\pi}$ as the momentum, energy and mass of the $N\pi$ system in the rest frame of $\Lambda_{b}^{0}$.
In the first line of Eq. (5), the CP asymmetries are the direct CPV generated by $|A|^{2}$, $|B|^{2}$, $|\tilde{A}|^{2}$ and $|\tilde{B}|^{2}$, as the squared modulus terms of $|P_{11}|^{2}$ and $|S_{11}|^{2}$ do not contribute to interference effects. For example,
$|A|^{2}-|\bar{A}|^{2}$ $\propto 2\mathcal{R}e(\lambda_{u}\lambda_{t}a_{1}a_{46+})-2\mathcal{R}e(\lambda^{*}_{u}\lambda^{*}_{t}a_{1}a_{46+})$ (7)
$\propto\sin(\Delta\phi_{w})\sin(\Delta\delta),$
where the weak phase difference $\Delta\phi_{w}$ stems from the CKM matrix elements $\lambda_{u}\lambda_{t}$ and $\lambda^{*}_{u}\lambda^{*}_{t}$, while the strong phase difference $\Delta\delta$ is from the effective Wilson coefficients $a_{1}$ and $a_{46+}$. CP violation in this case is induced by the interference between the tree and penguin amplitudes in the same partial wave, such as $A$ in the above formula.
The second line of Eq. (5) introduces a different CPV mechanism where the re-scattering strong phases of $S_{11}$ and $P_{11}$ make the contribution to the strong phase difference of CPV. The associated CP asymmetry, for example, is
$\mathcal{R}e\left[AP_{11}\tilde{A}^{*}S^{*}_{11}\right]-\mathcal{R}e\left[\bar{A}\bar{P}_{11}\bar{\tilde{A}}^{*}\bar{S}^{*}_{11}\right]$ (8)
$\propto\mathcal{R}e\big{[}(\lambda_{u}^{*}\lambda_{t}-\lambda_{u}\lambda_{t}^{*})(a_{46+}P_{11})(a_{1}^{*}S_{11}^{*})\big{]}$
$\hskip 11.38092pt+\mathcal{R}e\big{[}(\lambda_{u}\lambda_{t}^{*}-\lambda_{u}^{*}\lambda_{t})(a_{1}P_{11})(a_{46-}^{*}S_{11}^{*})\big{]}.$
The weak phase difference is from $\lambda_{u}^{*}\lambda_{t}$ and $\lambda_{u}\lambda_{t}^{*}$, whereas the strong phase differences originate from $a_{46+}P_{11}$ and $a_{1}^{*}S_{11}^{*}$, or $a_{1}P_{11}$ and $a_{46-}^{*}S_{11}^{*}$. The partial-wave amplitudes of $N\pi$ scatterings explicitly contribute to the strong phases. This type of CPV is induced by the interference between the tree and penguin amplitudes in different partial waves. For instance, the tree amplitude of the $1/2^{-}$ mode, $a_{1}S_{11}$, interferes with the penguin amplitude of the $1/2^{+}$ mode, $a_{46+}P_{11}$. Analogous to the interference between the tree amplitude of the $1/2^{+}$ mode, $a_{1}P_{11}$, and the penguin amplitude of the $1/2^{-}$ mode, $a_{46-}S_{11}$. Besides, since the $N\pi$ scattering amplitudes depend on the colliding energy, the corresponding CPV might vary as the changes of the invariant mass of the $N\pi$ system.
The using of the data of $N\pi$ scatterings is similar to the $\pi\pi\to K\bar{K}$ rescattering in $B$-meson three-body decays Bediaga:2013ela ; AlvarengaNogueira:2015wpj ; Cheng:2016shb ; Garrote:2022uub . The model-independent dispersive analysis of $\pi\pi\to K\bar{K}$ data Pelaez:2020gnd can lead to a very accurate description of the CP asymmetries of measured by LHCb Garrote:2022uub . However, the CPV mechanism with $N\pi$ scatterings is different from that of $B$-meson decays. In $B$-meson decays, the rescattering of $\pi\pi\to K\bar{K}$ happens between two channels, like $B^{\pm}\to\pi^{\pm}\pi^{+}\pi^{-}$ and $\pi^{\pm}K^{+}K^{-}$, with the weak phases in these two processes respectively. The $N\pi$ scatterings in $\Lambda_{b}$ decays contribute to one individual process but with different partial waves, which is more similar to the interference of $S$-wave and $P$-wave amplitudes in $B^{+}\to\pi^{+}\pi^{+}\pi^{-}$.
@@
In the end, it can be easily studied for the angular analysis on $\Lambda_{b}$ decays, benefited by the partial-wave amplitudes of $N\pi$ scatterings. The angular distributions are illustrated in Fig. 2 for $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The angle $\theta$ is defined between the directions of momentum of the final baryon $p(\Delta^{++})$ in the rest frame of $N\pi$ system and the momentum of $N\pi$ in the rest frame of $\Lambda_{b}^{0}$. The formula of angular distributions can be easily obtained by the helicity amplitudes.
The above CPV mechanism from $N\pi$ scatterings can also be applied in charmed baryon decays, such as $\Lambda_{c}^{+}\to(p\pi^{-})\pi^{+}$. Other scatterings can also generate similar CPV, for instance $N\pi\to\Lambda\bar{K}$ and $NK\to NK$. In this work, we will only discuss $\Lambda_{b}$ decays with $N\pi$ scattering for the introduction of this new CPV mechanism, and leave all the other discussions in the future.
The proposal of $N\pi$ scatterings can also pave the way for experimental analysis of heavy-flavor baryon decays. The amplitude analysis in experiments always suffer large uncertainties and difficulties from the too many baryon resonances. The using of data of $N\pi$ scatterings might help experimentalists do the partial-wave amplitude analysis.
III Numerical Discussions
Table 1: Global and regional CPV of $\Lambda_{b}$ decays stemming from $N\pi$ re-scatterings. $N\pi\to\Delta\pi$ contributes to the processes of $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$ and $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})\pi^{-}$, while $N\pi\to N\pi$ contributes to $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$ and $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$. The results are given in three scenarios of the form factors, with S1: $f_{1}=1.1$, $g_{1}=0.9$, S2: $f_{1}=g_{1}=1.0$, and S3: $f_{1}=0.9$, $g_{1}=1.1$. The predictions are provided in the invariant mass regions of $m_{N\pi}\in[1.2,1.9]$GeV for $N\pi\to\Delta\pi$, and $m_{N\pi}\in[1.1,2.5]$GeV for $N\pi\to N\pi$, which are constrained by the SAID data.
      decay processes        Scenarios         global CPV      CPV of $\cos\theta<0$       CPV of $\cos\theta>0$
$\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$ S1 5.9% 8.0% 3.6%
S2 5.8% 6.3% 5.3%
S3 5.6% 4.3% 7.0%
$\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})\pi^{-}$ S1 --4.1% --5.4% --2.4%
S2 --3.9% --3.9% --3.9%
S3 --3.6% --2.3% --5.3%
$\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$ S1 5.8% 8.2% 2.7%
S2 5.8% 8.0% 3.0%
S3 5.8% 7.8% 3.3%
$\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$ S1 --3.9% --3.9% --3.7%
S2 --3.9% --3.8% --4.3%
S3 --3.8% --3.6% --4.8%
Figure 3: CPV distributions of the Dalitz plots. They are the processes of $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$ and $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, from top to bottom. The results for each process are given in three scenarios of the form factors, with S1: $f_{1}=1.1$, $g_{1}=0.9$, S2: $f_{1}=g_{1}=1.0$, and S3: $f_{1}=0.9$, $g_{1}=1.1$, for the left, the middle and the right figures, respectively.
Based on the above CPV mechanism and the data of $N\pi$ scatterings, we can predict the global and regional CP asymmetries of $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The inputs of effective Wilson coefficients are given in Ali:1998eb . The inputs of $N\pi$ scattering amplitudes are model-independent and given in SAID . Since the data of scatterings of $N\pi\to N\pi$ and $\Delta\pi$ are very precise, we neglect the uncertainties of scattering amplitudes.
@@ III Numerical Discussions
Little is known about the form factors of $\Lambda_{b}\to N\pi$ at the current stage. But it is similar to the transition form factors of $B\to\pi\pi$ which have been studied by the heavy meson chiral perturbation theory Lee:1992ih ; Fajfer:1998yc ; Cheng:2007si ; Cheng:2020ipp and the light-cone QCD sum rules Cheng:2017smj ; Cheng:2017sfk . Therefore the form factors of $\Lambda_{b}\to N\pi$ can be calculated using the proven methods in principle. At the current stage, we will take some simple and typical values of these form factors, in order to reveal the new CPV mechanism. In the heavy quark limit, the form factors $f_{1}=g_{1}$. The values of $f_{1}$ and $g_{1}$ might be different by about $10\%$ when considering the corrections of heavy quark expansion Han:2022srw ; Manohar:2000dt . Besides, the absolute values of form factors are cancelled in the ratio of CPV definition. CPV relies only on the relative values of form factors. Without loss of generality, we naively take three sets of values, i.e. $f_{1}=g_{1}=1.0$ to give results as the central values, with those of $f_{1}=1.1$, $g_{1}=0.9$, and $f_{1}=0.9$, $g_{1}=1.1$ as the error estimations. In the following discussions, we will provide all the results of Table. 1 and Figs. 3 and 4 in three scenarios of the form factors, with S1: $f_{1}=1.1$, $g_{1}=0.9$, S2: $f_{1}=g_{1}=1.0$, and S3: $f_{1}=0.9$, $g_{1}=1.1$. This can be seen what are the results in the heavy quark symmetry, and how large the corrections beyond the heavy quark symmetry could affect. The mass regions of $p\pi^{0}$ and $\Delta^{++}\pi^{-}$ are taken as what SAID program provides, i.e. $m_{p\pi^{0}}\in[1.1,2.5]$GeV and $m_{\Delta^{++}\pi^{-}}\in[1.2,1.9]$GeV. Then we can get the numerical results using the above inputs.
Firstly, the global CPV in the mass regions of the $N\pi$ scattering system are given in Tables. 1. It can be found that the magnitudes of global CPV of these processes are around ±$\pm(4-6)\%$. Such values of CPV could be measured by LHCb for $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ and $(p\pi^{+}\pi^{-})\pi^{-}$ whose data samples are large enough with high detection efficiency of all the charged final-state particles. From Table. 1, the results of global CPVs do not change significantly with the varying values of form factors $f_{1}$ and $g_{1}$. The reason is that the global CPV is an overall and averaged effect, and thus is basically generated by the short-distance contributions where the strong phases are from the effective Wilson coefficients. The partial-wave amplitudes of the long-distance $N\pi$ scatterings usually affect the local CPVs in the Dalitz plots. Therefore, the predictions of the global CPVs can be used for larger regions to be compared with the experimental measurements. Even though the invariant mass regions are limited by the SAID data, the global CPV can be extended to larger regions and then compared to the LHCb measurements. The LHCb collaboration has measured the CPVs using the data sample of Run I as, $A_{CP}(\Lambda_{b}^{0}\to\Delta^{++}\pi^{-}\pi^{-})=(+0.1\pm 3.2\pm 0.6)\%$ and $A_{CP}(\Lambda_{b}^{0}\to\Delta^{++}\pi^{-}K^{-})=(+4.4\pm 2.6\pm 0.6)\%$ LHCb:2019jyj . Our predictions, $A_{CP}(\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})\pi^{-})\approx-4\%$ and $A_{CP}(\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-})\approx 5\%$, are consistent with the LHCb results within the uncertainties.
The opposite signs between the CPVs of emitted $K^{-}$ modes and $\pi^{-}$ modes are basically from the weak phases. The CKM matrix elements of emitted $K^{-}$ modes are $V_{ub}V_{us}^{*}$ for the tree contributions, and $V_{tb}V_{ts}^{*}$ for the penguin contributions. Therefore the CPVs are proportional to $\sin(\gamma)>0$ with $\gamma\sim 65^{\circ}$ ParticleDataGroup:2024cfk . For the emitted $\pi^{-}$ modes, the CKM matrix elements are $V_{ub}V_{ud}^{*}$ and $V_{tb}V_{td}^{*}$ in the tree and penguin contributions, respectively. The corresponding CPVs are proportional to $\sin(-\alpha)<0$ with $\alpha\sim 84^{\circ}$ ParticleDataGroup:2024cfk .
@@
The regional CPV in the Dalitz plots are even more interesting. Every partial-wave amplitude corresponds to different angular distribution. Then the interference effects in the different regions in the Dalitz plots would induce different CP asymmetries. The numerical results of regional CPV are given in Figure. 3. It could be found that the local CPV in some regions could reach the order of $10\%$, or be even as large as $30\%$. This is large enough for the observation in experiments. Besides, the form factors of $f_{1}$ and $g_{1}$ could affect the results significantly in some processes, which means the regional CPV are more sensitive to the corrections of heavy quark expansion, compared to the global CPV in Table. 1.
@@
It is also interesting that there exists large regions with CPV of positive or negative values in the Dalitz plot. This would be helpful for the experimental measurements. For example, it can be divided into two bins with $\cos\theta>0$ and $\cos\theta<0$ in all the four processes. The numerical results of CPV in each bin are given in Tables. 1.
@@
From the Dalitz plots, it could be clearly seen that the regional CPV significantly depend on $\cos\theta$  [^3] The invariant mass squares, $m_{p\pi^{0}}^{2}$ or $m_{\Delta^{++}\pi^{-}}^{2}$, related to the vertical axis in the Dalitz plots, are linearly proportional to $\cos\theta$. Taking $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$ as an example, $m^{2}_{\pi^{0}\pi^{-}}=$ $\left(\frac{m^{2}_{N\pi}+m^{2}_{\pi^{0}}-m^{2}_{p}}{2m_{N\pi}}+\frac{m^{2}_{\pi^{-}}+m^{2}_{\Lambda_{b}}-m^{2}_{N\pi}}{2m_{\Lambda_{b}}}\right)^{2}$ $-\Big{(}|\vec{p}_{\pi^{0}}|^{2}+|\vec{p}_{\pi^{-}}|^{2}+2|\vec{p}_{\pi^{0}}||\vec{p}_{\pi^{-}}|\cos\theta\Big{)}.$ Therefore, the regions of $\cos\theta>0$ and $\cos\theta<0$ are depicted as lower and upper zone in the Dalitz plot, respectively. . This is reasonable that the interference between different partial-wave amplitudes induce the dependence of CPV on $\cos\theta$. In order to see the dependence, it is very convenient to express the differential decay width on Legendre series $P_{n}(\cos\theta)$. It is known that the angular distributions are actually related to the Legendre polynomials. The Legendre distributions after integrating out scattering energy are
$\frac{d\Gamma}{d\cos\theta}\propto\sum_{n=0}\mathcal{L}_{n}P_{n}(\cos\theta)$ (9)
with $\mathcal{L}_{n}=(1,-0.10,0.20,-0.05,0.009,0.05)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, $(1,-0.39,0.24,-0.08,-0.09,0.05)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$, (1,0.4,0.4,0.5,0.03,0.12,(1,-0.4,0.4,-0.5,-0.03,-0.12,( 1 , - 0.4 , 0.4 , - 0.5 , - 0.03 , - 0.12 , 0.005)-0.005)- 0.005 ) for $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$, and (1,1.25,1.1,0.7,(1,-1.25,1.1,-0.7,( 1 , - 1.25 , 1.1 , - 0.7 , 0.2,0.1,0.01)0.2,-0.1,0.01)0.2 , - 0.1 , 0.01 ) for $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, respectively, in the case of $f_{1}=1.1$ and $g_{1}=0.9$. It can be found that the first few moments make large contributions, with small contributions from higher moments. This can be understood that the higher partial-wave amplitudes are suppressed, while the lower partial-wave amplitudes make the dominant contributions.
Figure 4: CP asymmetries of the Legendre moments in the unit of percent, with the horizontal axis $n=1,2,\cdots$, of $P_{n}(\cos\theta)$. They are the processes of $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$ and $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, from top to bottom. The results for each process are given in three scenarios of the form factors, with S1: $f_{1}=1.1$, $g_{1}=0.9$, S2: $f_{1}=g_{1}=1.0$, and S3: $f_{1}=0.9$, $g_{1}=1.1$, for the left, the middle and the right figures, respectively.
@@
Then the large regional CPV could also be measured by the Legendre moments. It can be easily realized by the experiments using the differential width convoluted with the Legendre polynomials. The results of CPV of Legendre moments are given in Figure. 4. The interesting result is that the CPV of the first moment $\cos\theta$ could be as large as ±$\pm 70\%$ or $40\%$ in $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$, which can be easily measured and tested in experiments.
@@
In the end, we strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ for the first observation of baryon CPV. $\Lambda_{b}^{0}\to p\pi^{+}\pi^{-}K^{-}$ has the largest data sample of charmless $b$-baryon decays at LHCb LHCb:2019jyj . It could be measured in three ways. Firstly, the global CPV with the value of $6\%$ could be reached with all the data sample. Secondly, the CPV in the Dalitz plot can be measured with the regional CPV of $(10-20)\%$. Thirdly, it is also realizable to measure the CPV of moment of $\cos\theta$ with a value as large as ±$\pm 70\%$ or $40\%$. Using the data samples of integrated luminosity of 3fb-1 at LHCb Run 1, the total CPV of $\Lambda_{b}^{0}\to p\pi^{+}\pi^{-}K^{-}$ is measured to be $(+3.2\pm 1.1\pm 0.6)\%$, while that of $\Lambda_{b}^{0}\to\Delta^{++}\pi^{-}K^{-}$ is $(+4.4\pm 2.6\pm 0.6)\%$ LHCb:2019jyj . The current total data with 9fb-1 collected by LHCb Run 1 and 2 would be larger by around five times due to the larger production rate at Run 2. So the precision of the global CPV in $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ could be around $1\%$ YanxiZhang . LHCb Run 3 is now collecting even more data during 2024 and 2025, with the total integrated luminosity up to 23fb-1. Therefore, it can be expected that the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ could be observed in the very near future, with the global CPV of around $(5-6)\%$ and the experimental precision lower than $1\%$.
IV Conclusions